OpenStudy (anonymous):

How to Find out; the Resultant Displacement & the Direction of the resultant (with respect to +ve x-axix) If a Car travels 200 km due EAST and then 150 km 60 (degrees) North Of East.

OpenStudy (turingtest):

Draw a vector diagram:it looks Law of cosines will take care of this in a flash\[\large d^2=a^2+b^2-2ab\cos D \]where D is the angle that subtends side d (which is the displacement, and a and b are the other sides of the triangle